OpenStudy (anonymous):

-6x-y=-25 -3x+4y=19 Solve Eacg System By Elimination Steps Only Please

OpenStudy (lynfran):

mutiple the 2nd equation throughout by 2 to get .equation 3\[-6x+8y=38\]..then take away the 1st and 3rd equatio

OpenStudy (anonymous):

you could also multiply the first equation by 4 and then add the 2 equations

OpenStudy (anonymous):

Can Someone Show Me The Steps Because I Got The Answer Already

OpenStudy (lynfran):

\[-6x-y=-25\]\[-(-6x+8y=38)\]

OpenStudy (lynfran):

note that the x will cancel leaving \[-9y=-63\]then we can solve for y by dividing both side of the equation by -9

OpenStudy (lynfran):

so what will y=?

OpenStudy (anonymous):

-72

OpenStudy (lynfran):

\[y=\frac{ -63 }{ -9 }\]

OpenStudy (anonymous):

Oh So Dont Solve It ?

OpenStudy (lynfran):